egb520:

The volumes of two cylindrical cans of the same shape vary directly as the cubes of their radii. If a can with a six-inch radius holds 1½ pints, how many gallons will a similar can with a 24-inch radius hold?

AngeI:

Hello welcome to QuestionCove ;3 Sorry Im still gathering my thoughts on this, please give me a few mins

egb520:

Ofc! take your time. Thank you!

AngeI:

So 1 1/2 pints= 3/16th of a gallon I do believe that we would use R^3 for the sizes 6^3*x=24^3*3/16 216*x=13824*3/16 216*x=2592 Then you would divide both sides by 216 getting x=12 gallons The numbers get rather big but thats just because of the "^3" You'd normally be able to use a caculator on these problems so do not worry c:

egb520:

Thank you! I appreciate it a lot!

AngeI:

Of course, I hope you enjoy your time on this site c: